OpenStudy (anonymous):

Absolute value equation. Solve for x: -3|2x + 6| = -12

OpenStudy (abbles):

x = -1

OpenStudy (abbles):

Any questions?

OpenStudy (anonymous):

Thank you! And well I needed two answers, one positive answer and one negative.

OpenStudy (abbles):

Hm... I don't see any other way to solve it. Maybe -1 and 1 as your answers?

OpenStudy (anonymous):

I don't have those for my answer choices. These are my answer choices: a. x = 1 and x = 5 b. x = -1 and x = -5 c. x = -9 and x = 3 d. No solution But I do know that after you solve 2x+6= 5 -6 -6 2x=1 But then after that you have to write the quation like this 2x+6= -5 And I tried to solve it and got -11 as my answer but thats not a choice so I know it's wrong so I'm stuck

OpenStudy (abbles):

Maybe one of the answers is an extraneous solution, meaning it doesn't work out. Therefore, when plugged back into the equation it doesn't solve. That would make sense... I think it's b because it is the only answer that includes -1. Either b or no solution.

OpenStudy (anonymous):

Alright thank you I will go with that one and see what I get!

OpenStudy (abbles):

Okay, let me know what the answer is :)

OpenStudy (anonymous):

Will do(:

OpenStudy (anonymous):

It was correct! Lol Thank you(:

OpenStudy (abbles):

Haha I'm glad :) welcome